Base/Acid Ratios in Buffers

Just as pH is the negative logarithm of [H3O+], pKa is the negative logarithm of Ka,

pKa=?logKa

The Henderson-Hasselbalch equation is used to calculate the pH of buffer solutions:

pH=pKa+log[base][acid]

Notice that the pH of a buffer has a value close to the pKa of the acid, differing only by the logarithm of the concentration ratio [base]/[acid].

Part A

Acetic acid has a Ka of 1.8
Answer #1

Part A. We will use the Henderson's -Hasselbalch equation,

pH = pKa + log(A-/HA)

given here is Ka = 1.8 x 10^-5

thus, pKa = -log(1.8 x 10^-5)

                = 4.744

A. pH = 4.744 + log 1 / 10

          = 4.744 -1

          = 3.744

B. pH = 4.744 + log 10 / 1

          = 4.744 + 1

          = 5.744

C. pH = 4.744 + log 1 / 1

          = 4.744

Part B. Again we will use the Henderson-Hasselbalch equation,

pH = pKa + log([A^-]/[HA])

The reaction equation will be,

NH3(aq) + H2O ? [NH4]^+ + OH^-

So we have the equation,

pH = pKa + log([NH3]/[NH4^+]

Given, Kb = 1.8